# Measurement of the work statistics of an open quantum system using a quantum computer

### v1

@@ Abstract
We report on the experimental measurement of the work statistics of a genuinely open quantum system using a quantum computer. Such measurement has remained elusive thus far due to the inherent difficulty in measuring the total energy change of a system-bath compound (which is the work) in the open quantum system scenario. We overcome this difficulty by extending the interferometric scheme, originally conceived for closed systems, to the open system case and implement it on a superconducting quantum computer, taking advantage of the relatively high levels of noise on current quantum hardware to realize an open quantum system. We demonstrate that the method can be used as a diagnostic tool to probe physical properties of the system-bath compound, such as its temperature and specific transitions frequencies in its spectrum. Our experiments corroborate that the interferometric scheme is a promising tool to achieve the long-sought experimental validation of the Jarzynski equality for arbitrary open quantum systems.
@@ I Introduction
In order to see how the method can be extended to the open quantum system case, consider, without loss of generality, the prototypical case when $V(t)=\lambda(t)Q$ (with $\lambda$ a real function such that $\lambda(t)=0$ for $t\notin(0,\tau)$ and $Q$ a system operator). In this case, the interferometric scheme is implemented by the evolution generated by the following System-Ancilla Hamiltonian in the time interval $t\in(0,\tau+u)$
@@
$f_{1}(t)=\lambda(t)\xi_{[0,\tau]}(t),\quad f_{0}(t)=\lambda(t-u)\xi_{[u,\tau+u]}(t)\,,$ (10)
with $\xi_{[a,b]}(t)$ denoting the indicator function of the interval $[a,b]$.
If the system that is manipulated by the action of the external force $Q$ is open (i.e., interacting with a bath), then the Hamiltonian that properly describes the $SBA$ dynamics (SB system plus ancilla), assuming the ancilla does not interact with the bath, becomes
$H_{SBA}(t)=H_{0}^{SB}\otimes\mathds{1}_{A}+f_{1}(t)Q^{SB}\otimes\Pi_{1}+f_{0}(t)Q^{SB}\otimes\Pi_{0}\,,$ (11)
where $Q^{SB}=Q\otimes\mathds{1}_{B}$, and
@@
Upon inspection of Eq. (11), we recognize the crucial fact that such a Hamiltonian generates a quantum circuit (see Fig. 1d) that encodes in the ancilla state, via Eq. (4), the function $g(u)$ associated to the the driven open quantum system
$H_{SB}(t)=H_{0}^{SB}+\lambda(t)Q^{SB}\,.$ (13)
If the initial SB preparation commutes with $H_{0}^{SB}$, then $g(u)$ is indeed the work characteristic function $\chi(u)$ of the driven open system, and its Fourier transform gives the corresponding TPM work PDF, accounting for the energy changes $w=E_{m}^{SB}-E_{n}^{SB}$ occurring in the whole SB system ($E_{m}^{SB}$ denoting the eigenvalues of $H^{0}_{SB}$).
@@
In our experiments, the resonant frequency of the qubit is $\omega=4.8$ GHz, which sets $\hbar\omega=3.2$ $\mu$eV. Due to the openness of the system qubit in our experiment, and other possible sources of error, we expect to observe a deviation in our measured work PDF from the ideal work distribution $p^{id}(w)$.
@@
First, note that the closed-system work PDF does not exactly match the ideal work PDF $p^{id}(w)$, Eq. (23), despite the fact that it was produced with a noise-free quantum emulator. This discrepancy is due to the finite sampling range and rate of $g(u)$. Note from Eq. (14) that in order to exactly recover the quantity $q_{>}(w)$, one needs to sample $g(u)$ on the continuum infinite set $[0,\infty)$. Due to constraints related to access and compute time on the IBM quantum computers, we were limited to sampling 900 values for $u$. To achieve sufficient resolution and range of work values in the resulting work PDF, a step-size of $\Delta u=0.08$ between sampled values of $u$ was chosen, resulting in a sampling range of $[0,72]$ for $u$ in units of $[\mu\text{eV}]^{-1}$. The same sampling rate and range were used for all work PDF plots.
@@
Figure 2: Work PDFs measured via the interferometric scheme on real quantum computer versus a quantum emulator. (a) Comparison of work PDF measured from the quantum computer (blue dotted curve) versus the work PDF computed with a noise-free quantum emulator (solid gray curve) with the qubit initialized in the pure ground state. Inset: Measurement of $2\mathbb{RE}[q_{>}(w)]$ on the quantum computer when the system qubit is intentionally initialized in a coherent state. (b) Comparison of work PDF measured from the quantum computer (blue dotted curve) versus the work PDF computed with a noise-free quantum emulator (solid red curve) with the qubit initialized in a thermal state at the measured effective initial temperature $T_{0}=10.7$ mK.
@@
[…]
@@ IV Towards the verification of the Jarzynski Equality
We are, however, able to establish whether or not our experimental data are consistent with the validity of the Jarzynski equality. To do so, we performed a second experiment involving two sets of circuits: one set identical to the ones described above, and the other set with the the $S$ qubit intentionally prepared in its excited state, which simply requires enacting an $X$-gate on the qubit. Just as preparation of the ground state occurred with some finite fidelity (associated to the finite effective temperature $T_{0}$), preparation of the excited state likewise occurs with a finite fidelity, which can be associated with a negative temperature $T_{1}$. Using calibration circuits consisting solely of measurement, we measured the effective initial temperature of the qubit when it was prepared in the ground state to be $T_{0}=8.5$ mK, and analogously measured the effective initial temperature of the qubit when it was prepared in the excited state to be $T_{1}=-8.8$ mK.
We denote with $p(w,T_{1})$ the work PDF measured with the qubit initialized at an effective temperature $T_{1}$, and analogously, we denote with $p(w,T_{0})$ the work PDF measured with the qubit initialized at temperature $T_{0}$. Mixing the two work PDFs via a convex linear combination with weights $q$ and $1-q$,
$p(w,T)=qp(w,T_{0})+(1-q)p(w,T_{1})$ (28)
@@
$T=\frac{\hbar\omega}{k_{B}}\left[\ln\frac{qf(\hbar\omega\beta_{0})+(1-q)f(\hbar\omega\beta_{1})}{qf(-\hbar\omega\beta_{0})+(1-q)f(-\hbar\omega\beta_{1})}\right]^{-1}$ (29)
@@
$T_{B}\sim 39\text{ mK}$ (31)
denoted by the gray vertical line in Figure 3. While we do not expect this estimate to be extremely precise due to the effects of numerous sources of error (e.g., discrete Fourier transform over truncated range, quantum computational error, shot noise, etc.), this temperature nevertheless sits consistently between the nominal temperature of the cryostat housing the quantum processor of $15$ mK and independent measurements of qubit temperatures on the IBM quantum processors reported in Ref. [24], ranging between $\sim 40-60$ mK. Said in different terms, our data are consistent with the validity of the Jarzynski equality and thus constitute a promising method to achieve a fully fledged verification thereof, for open quantum systems.
@@ V Discussion
Last but not least, our experiments demonstrate that the interferometric scheme is a promising method for reaching the long-sought experimental verification of the Jarzynski equality for arbitrary open quantum systems [11]. While we cannot claim to have achieved such a verification, as we lack a means of independent measurement of the temperature of the bath, our data are nevertheless consistent with the Jarzynski equality. This indicates that a dedicated experiment with finer controls than those available to the general user of a quantum computer could very likely reach the goal.
Using a reversed but equivalent logic, knowing that the measured work PDF of an open quantum system indeed obeys the Jarzynski equality allows one use the work PDF to estimate the temperature of the bath. Seen from this perspective, the interferometric scheme could be used as a thermometer. In our case, due to the error affecting the data, such an estimate of the bath temperature was very rough, although consistent with other existing estimates. Once again, with a tailor made experiment one could achieve a much higher thermometric precision.

### v2

@@ Abstract
We report on the experimental measurement of the work statistics of a genuinely open quantum system using a quantum computer. Such measurement has remained elusive thus far due to the inherent difficulty in measuring the total energy change of a system-bath compound (which is the work) in the open quantum system scenario. We overcome this difficulty by applying the interferometric scheme, originally conceived for closed systems, to the open system case and implement it on a superconducting quantum computer, taking advantage of the relatively high levels of noise on current quantum hardware to realize an open quantum system. We demonstrate that the method can be used as a diagnostic tool to probe physical properties of the system-bath compound, such as its temperature and specific transition frequencies in its spectrum. Our experiments corroborate that the interferometric scheme is a promising tool to achieve the long-sought experimental validation of the Jarzynski equality for arbitrary open quantum systems.
@@ II.1 Closed case
In order to see how the method applies to the open quantum system case, consider, without loss of generality, the prototypical case when $V(t)=\lambda(t)Q$ (with $\lambda$ a real function such that $\lambda(t)=0$ for $t\notin(0,\tau)$ and $Q$ a system operator). As thoroughly illustrated in Ref. [13], in the closed case, the interferometric scheme is implemented by the evolution generated by the following System-Ancilla Hamiltonian in the time interval $t\in(0,\tau+u)$
@@
$\begin{split}f_{1}(t)=\begin{cases}\lambda(t),&0\leq t<\tau\\ 0,&\tau\leq t\leq\tau+u\\ \end{cases}\\ f_{0}(t)=\begin{cases}0,&0\leq t<u\\ \lambda(t-u),&u\leq t\leq\tau+u\\ \end{cases}\end{split}$ (10)
Basically, the method consists in running the protocol $\lambda(t)$ for a time $\tau$ followed by a delay time $u$ during which $\lambda$ is kept at zero, conditioned on the ancilla being in state 00; and keeping $\lambda$ at zero for a time $u$ followed by the application of $\lambda(t)$ for a time $\tau$, conditioned on the ancilla being in state $1$. Thus $u$ is, physically, the delay time between the application of the driving $\lambda(t)$ in the two subspaces of the ancilla, which is responsible of the interference pattern collected in the state of the ancilla itself. Accordingly $\tau+u$ is the total time duration of the interferometric circuit. As shown in Refs. [13, 14], the Hamiltonian (9) imprints in the ancilla state, the function $g(u)$ associated with the driven closed quantum system $H_{S}(t)=H_{0}+\lambda Q$.
II.2 Open case
If you apply the very same driving described above but your system S is open (i.e., interacting with a bath), then the Hamiltonian that properly describes the $SBA$ dynamics (SB system plus ancilla), assuming the ancilla does not interact with the bath, is
$H_{SBA}(t)$ $=$ (11)
$H_{0}^{SB}\otimes\mathds{1}_{A}+f_{1}(t)Q\otimes\mathds{1}_{B}\otimes\Pi_{1}+f_{0}(t)Q\otimes\mathds{1}_{B}\otimes\Pi_{0}$
where
@@ II.2 Open case
Upon inspection of Eq. (11), we recognize fact (and this is the most crucial point of our whole analysis) that such a Hamiltonian generates a quantum circuit (see Fig. 1d) that encodes in the ancilla state, via Eq. (4), the function $g(u)$ associated to the the driven open quantum system
$H_{SB}(t)=H_{0}^{SB}+\lambda(t)Q\otimes\mathds{1}_{B}\,.$ (13)
In other words, by enacting one and the same driving on the $SA$ compound, namely the $SA$ driving $f_{1}(t)Q\otimes\Pi_{1}+f_{0}(t)Q\otimes\Pi_{0}$ , one gets for free either the closed-S system work PDF or the open-S system work PDF, depending on whether the S-system is closed or open, respectively [17]. More precisely, if the initial SB preparation commutes with $H_{0}^{SB}$, then $g(u)$ is indeed the work characteristic function $\chi(u)$ of the driven open system, and its Fourier transform gives the corresponding TPM work PDF, accounting for the energy changes $w=E_{m}^{SB}-E_{n}^{SB}$ occurring in the whole SB system ($E_{m}^{SB}$ denoting the eigenvalues of $H^{0}_{SB}$).
II.3 Dealing with the constraint $u\geq 0$
@@
In our experiments, the resonant frequency of the qubit is $f=4.85$ GHz (where $f=\omega/2\pi$), which sets $hf=\hbar\omega=20.04$ $\mu$eV. Due to the openness of the system qubit in our experiment, and other possible sources of error, we expect to observe a deviation in our measured work PDF from the ideal work distribution $p^{id}(w)$.
@@
First, note that the closed-system work PDF does not exactly match the ideal work PDF $p^{id}(w)$, Eq. (23), despite the fact that it was produced with a noise-free quantum emulator. This discrepancy is due to the finite sampling range and rate of $g(u)$. Note from Eq. (14) that in order to exactly recover the quantity $q_{>}(w)$, one needs to sample $g(u)$ on the continuum infinite set $[0,\infty)$. Due to constraints related to access and compute time on the IBM quantum computers, we were limited to sampling 900 values for $u$. To achieve sufficient resolution and range of work values in the resulting work PDF, a step-size of $\Delta u=0.013$ between sampled values of $u$ was chosen, resulting in a sampling range of $[0,11.5]$ for $u$ in units of $[\mu\text{eV}]^{-1}$. The same sampling rate and range were used for all work PDF plots.
@@
Figure 2: Work PDFs measured via the interferometric scheme on real quantum computer versus a quantum emulator. (a) Comparison of work PDF measured from the quantum computer (blue dotted curve) versus the work PDF computed with a noise-free quantum emulator (solid gray curve) with the qubit initialized in the pure ground state. Inset: Measurement of $2\mathbb{RE}[q_{>}(w)]$ on the quantum computer when the system qubit is intentionally initialized in a coherent state. (b) Comparison of work PDF measured from the quantum computer (blue dotted curve) versus the work PDF computed with a noise-free quantum emulator (solid red curve) with the qubit initialized in a thermal state at the measured effective initial temperature $T_{0}=67$ mK.
@@
To investigate the origin of the third peak, we examined the fidelity of the initial preparation of the $S$ qubit in the ground state. To do so, we executed a calibration circuit comprising only a measurement of the $S$ qubit. Across 1024 runs of the calibration circuit, the qubit was measured in the ground state with a probability of $p_{0}\simeq 0.97$. This would imply an effective initial temperature of $T_{0}=67$ mK for the qubit. To ascertain how this initial temperature will contribute to the third peak, we performed a noise-free emulation of the circuit in Fig. 1b with the $S$ qubit initialized in a thermal state at this temperature $T_{0}$.
@@ IV Towards the verification of the Jarzynski Equality
We are, however, able to establish whether or not our experimental data are consistent with the validity of the Jarzynski equality. To do so, we performed a second experiment involving two sets of circuits: one set identical to the ones described above, and the other set with the the $S$ qubit intentionally prepared in its excited state, which simply requires enacting an $X$-gate on the qubit. Just as preparation of the ground state occurred with some finite fidelity (associated to the finite effective temperature $T_{0}$), preparation of the excited state likewise occurs with a finite fidelity, which can be associated with a negative temperature $T_{1}$. Using calibration circuits consisting solely of measurement, we measured the effective initial temperature of the qubit when it was prepared in the ground state to be $T_{0}=83$ mK, and analogously measured the effective initial temperature of the qubit when it was prepared in the excited state to be $T_{1}=-87$ mK.
We denote with $p(w,T_{1})$ the work PDF measured with the qubit initialized at an effective temperature $T_{1}$, and analogously, we denote with $p(w,T_{0})$ the work PDF measured with the qubit initialized at temperature $T_{0}$. Mixing the two work PDFs via a convex linear combination with weights $r$ and $1-r$,
$p(w,T)=rp(w,T_{0})+(1-r)p(w,T_{1})$ (28)
@@
$T=\frac{\hbar\omega}{k_{B}}\left[\ln\frac{rf(\hbar\omega\beta_{0})+(1-r)f(\hbar\omega\beta_{1})}{rf(-\hbar\omega\beta_{0})+(1-r)f(-\hbar\omega\beta_{1})}\right]^{-1}$ (29)
@@
$T_{B}\sim 290\text{ mK}$ (31)
denoted by the gray vertical line in Figure 3. This estimate appears higher than independent measurements of qubit temperatures on the IBM quantum processors reported in Ref. [24], ranging between $\sim 40-60$ mK. However, it is a rather rough estimation due to the effects of numerous sources of error (e.g., discrete Fourier transform over truncated range, quantum computational error, shot noise, etc.). Nevertheless, we demonstrate how our method could potentially be used to estimate the temperature of qubits and their environment and propose that a bespoke experimental device with fine-tuned control over all qubit parameters, however, would likely succeed in obtaining a more accurate measurement of the temperature.
@@ V Discussion
Last but not least, our experiments demonstrate that the interferometric scheme is a promising method for reaching the long-sought experimental verification of the Jarzynski equality for arbitrary open quantum systems [11]. While we cannot claim to have achieved such a verification, as we lack a means of independent measurement of the temperature of the bath, our data are nevertheless not inconsistent with the Jarzynski equality. This indicates that a dedicated experiment with finer controls than those available to the general user of a quantum computer could very likely reach the goal.
Using a reversed but equivalent logic, knowing that the measured work PDF of an open quantum system indeed obeys the Jarzynski equality allows one use the work PDF to estimate the temperature of the bath. Seen from this perspective, the interferometric scheme could be used as a thermometer. In our case, due to the error affecting the data, such an estimate of the bath temperature was very rough. Once again, with a tailor made experiment one could achieve a much higher thermometric precision.